OpenStudy (anonymous):

Simplifying??? (2/3r^2s^3z^6)

OpenStudy (anonymous):

Is your question like this \[\frac{2}{3r^2s^3z^6}\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok, pls wait

OpenStudy (anonymous):

\[\frac{2}{3r^2s^3z^6} =\frac{2r^{-2}s^{-3}z^{-6}}{3} ={2 \times 3^{-1} r^{-2}s^{-3}z^{-6}}\]